# While recognizing actions in videos, LMMs struggle to detect core interaction events

### v1

@@ Abstract
Large multi-modal models (LMMs) show increasing performance in realistic visual tasks for images and, more recently, for videos. For example, given a video sequence, such models are able to describe in detail objects, the surroundings and dynamic actions. In this study, we explored the extent to which these models ground their semantic understanding in the actual visual input. Specifically, given sequences of hands interacting with objects, we asked models when and where the interaction begins or ends. For this purpose, we introduce a first of its kind, large-scale dataset with more than 20K annotated interactions on videos from the Something-Something-V2 dataset. 250 AMTurk human annotators labeled core interaction events, particularly when and where objects and agents become attached (’contact’) or detached (’release’). We asked two LMMs (Qwen-2.5VL and GPT-4o) to locate these events in short videos, each with a single event. The results show that although the models can reliably name the target objects, identify the action and provide coherent reasoning, they consistently fail to identify the frame where the interaction begins or ends and cannot localize the event within the scene. Our findings suggest that in struggling to pinpoint the moment and location of physical contact that defines the interaction, the models lack the perceptual grounding required for deeper understanding of dynamic scenes.
@@ 1 Introduction
Discovering and understanding actions and interactions are fundamental cognitive capabilities of humans and other intelligent beings, necessary in interpreting and planning dynamic events between objects and agents in the surrounding environment [radvansky2014event]. Infants develop early sensitivity to spatiotemporal continuity in simple events, for example, the physical contact between an agent and a target object. This sensitivity to the causality of perception guides infants at a very young age, in learning to detect and interpret interactions between objects and agents, including launching, entraining and expulsion events [michotte1963perception, leslie1982causality, saxe2006perception, baillargeon1995physical].
@@
Computationally, recent vision models showed increasing performance in recognizing actions and interactions in realistic video sequences [morshed_human_action_survey2023, gong_action_survey2025, tang_llm_video_survey2025]. Some models include special architectural designs to improve on the internal representation learning [long_videoprism2024], while others use a common architecture, but train on very large unlabeled video datasets utilizing self-supervised learning paradigms [tong2022videomae, liu_vlm2023, ishan_videossl2024]. The introduction of large multi-modal models (LMMs), allowed the combination of semantic information from large language models (LLMs) and foundational visual representations, thus allowing to generalize to unseen videos and actions without explicit training [tang_llm_video_survey2025].
Despite the increasing success of models in generalizing to more complex tasks without explicit training, recent studies revealed fundamental limitations in the models ability to reason about the performed tasks and develop human-like generalizable understanding [cherian_smart2023, buschoff_nature2025, hou2025do]. We are interested to explore whether the enhanced generalization behavior in interpreting actions in LMMs, reflects a deeper understanding of dynamic events at the core of interactions, similar to intelligent beings, or does it merely reflect a seemingly convincing ”story telling” ability of reliably detected objects in the scenes?
@@
S-Figure 1: Collecting human annotations for interactions using Amazon Mechanical Turk platform. Human subjects were asked to annotate core interaction events in videos from SSv2 dataset [goyal2017something]. Shown here are example annotations for ’contact’ and ’release’ events, where the target object (white candle) comes in contact with a hand (left) and a surface (middle), or is detached from the hand (right). The annotations include the event type, the kind of agent-object pair and the spatiotemporal location of the event (frame and image coordinates).
@@
For this purpose we introduce a large-scale dataset – The Contact-Release Interaction Dataset. This first of its kind dataset consists of more than 20K annotated interactions, based on $10,000$ action videos from the ”Something-Something v.2” (SSv2) dataset [goyal2017something]. Using the Amazon Mechanical Turk (AMTurk) crowd sourcing platform, we conducted a survey, in which 250 human annotators labeled $24,222$ core interaction events, including: (i) the type of agent acting upon the target object (e.g., a hand or another object), (ii) the type of core interaction event, i.e, ’contact’ if a target object becomes attached to an agent, or ’release’ if an object becomes detached from an agent, (iii) the spatiotemporal location of the event, i.e., frame number and image coordinates (see Fig. 1 and Sec. 3).
@@
Using the new annotations, we conducted a series of experiments to evaluate the ability of current LMMs to detect the spatio-temporal location of core interaction events in real-world video sequences. The experiments were conducted under several In-Context-Learning regimes, while applying two modifying conditions on the models’ prompts – Reasoning and Grounding – inspired by earlier studies [shao_visual_cot2024, wei_cot_prompting2022, guo_deepseekr12025]. In the evaluation, we used an open source version of Alibaba’s Qwen-2.5VL-72B [Qwen2.5-VL] and OpenAI’s online GPT-4o model [openai2024gpt4o].
@@
  • Introduce a large scale dataset, based on 10K action videos from SSv2 dataset, with more than 20K first of their kind human annotations of core interaction events (’Contact’ and ’Release’). The annotations include details on the event and agent types, as well as the spatiotemporal locations of the events.
  • A set of prompting experiments under several In-Context-Learning regimes and modifying conditions, including Reasoning and Grounding.
  • A discussion around the question: Does the enhanced generalization behavior of LMMs in action recognition reflect an improved video understanding, or is it merely a convincing ’story telling’ about detected objects in visual scenes?
@@ 2 Related Work
S-Figure 2: A schematic flow chart of the experiments under the different In-Context-Learning (ICL) regimes (i.e., ZS, OS, TS) and modulating conditions. The blocks represent different components of intermediate procedures. Each row represents an experiment using a particular ICL regime and condition (the experiment flow is directed left to right). The CNTX block indicates an introductory prompt about the agent. The EXMP block represents a prompt of an example, including the task instruction, an input video and the correct response for this example. The RSN block indicates a prompt instructing the model to include in the response a step-by-step description of the reasoning behind the predicted answer. The GRND block represents a prompt instructing the model to describe the content of the input video and the instructing prompt. In this block, the model provides an intermediate response, prior to the main task. The TST block indicates the prompt of the main test task, including the instruction and test video (see Sec. 4 for more details).
Video understanding was thoroughly researched in the past due to its high value to the advancement in the domain of AI. Unlike action recognition in video, such as identifying people jumping, playing tennis etc., video understanding is a more complex task, often requiring a high level of generalization. Recent studies in this area of research, such as Maaz et al. [maaz2024video], showed that an a ChatGPT agent can successfully answer complex questions when prompted with image data together with the verbal question. Wu et al. [wu2025f] used segmentation masks, which the model provided in a grounding step, to further improve the model’s understanding of the input images. Shao et al. [shao_visual_cot2024] showed that by asking the model to produce a Chain-of-Thought of a general task related to the image of the main task, improves significantly the final answer of the LMM. Tian and Wu [tian2025llm] also used a prompt tuning technique to improve action recognition performance of an LMM agent, but their method requires training of additional adapter models that embed the prompt, actions and the image to be tockenized and sent to the LMM. Chen et al. [chen2025visrl] employed a two-step reinforcement learning (RL) technique, where in the fist step the LMM agent is prompted to extract a region of interest (ROI) from an image based on the query; then this image crop, together with the actual test image and the associated prompt, are fed into the LMM for the actual answer. They also used the correctness of the prediction of the model on the test task as a guiding signal for models’ improvement. Qi et al. conduct a comprehensive prompt-probing study revealing fundamental limitations in multimodal LLMs’ visual grounding capabilities [Shuhan_llm_limit2023]. Their analysis, which is focused exclusively on image-based task, shows that these models often rely more on textual priors than on actual visual evidence, and that their visual understanding remains shallow despite producing coherent textual explanations. In contrast, our study investigates grounding in video settings, where models must additionally reason over temporal dynamics, object–object interactions, and event boundaries. This highlights a complementary and more challenging dimension of multi-modal grounding.
@@ 3 Dataset
S-Table 1: New annotated dataset of core interaction events. Videos and corresponding labeled action templates were taken from the Something-Something v.2 dataset [goyal2017something].
@@
  ’Contact’ events   $13816$
  ’Release’ events   $10406$
  ’Hand-object’ interactions   $12550$
  ’Object-object’ interactions   $5653$
  ’Object-surface’ interactions   $6019$
In this section we describe the dataset used in our experiments. As we explore the understanding of core interaction elements, for example the moment a hand picks up a target object including the spatial location of the hand and object in the scene, we introduce here a new large scale dataset – The Contact-Release Interaction Dataset. This first of its kind dataset consists of more than 20K annotated interactions on $10,130$ action videos from the ”Something-Something-V2” (SSv2) dataset [goyal2017something]. The SSv2 dataset is a collection of more than 200K labeled video clips of humans performing various everyday actions with objects in natural settings. The original dataset labels describe actions as generic templates, for example, ”putting something into something”, ”turning something upside down” and ”covering something with something”. The labels also include the actual object names in the template placeholder (”something”) for each video.
In this work we utilized the AMTurk platform to conduct a survey, in which 250 human annotators labeled $24,222$ core interaction events, including: (i) the type of agent acting upon the target object (e.g., a hand or another object), (ii) the type of core interaction event, i.e, ’contact’ if a target object becomes attached to an agent, or ’release’ if an object becomes detached from an agent, (iii) the spatiotemporal location of the event, i.e., frame number and image coordinates (see Fig. 1). The survey and the annotations collection procedure were approved by the institutional review board of the Weizmann Institute of Science, Rehovot, Israel. All human subjects gave informed consent before participating in the survey. Details about the new annotated dataset, including the videos and the labels are summarized in Tab. 1. The dataset is publicly available at ssv2-contact-release-interaction-dataset.
@@ 4 Experimental Design
In this section we describe in more detail the experiments we conducted in the course of the presented study. We evaluated LMMs under three In-Context-Learning (ICL) [dong_icl_survey2024] regimes: zero-shot, one-shot and two-shot.
We are interested in the detection of the spatiotemporal location of core interaction events in video sequences. In our experiments, we focused on the detection of the frame in the video, where a core interaction (’contact’ or ’release’) event occurs. Since the original videos in the SSv2 dataset may contain several core events, we extracted 3 isolated events for a subset of 33 videos, resulting with our experimental dataset, consisting of 99 short sequences, each with 10 frames consisting of a single core event.
To better understand the models behavior, we employed two explainability methods, that were used in previous studies with LMMs, as modulating conditions: (i) Grounding, (ii) Reasoning. In our experiments we evaluated their relative influence on the models performance. Fig. 2 presents the different experimental settings, including the ICL regimes and modulating conditions.
@@
In the OS regime, models are provided with a single example of the main task prior to performing the task on the test video sequence. The example, includes a video with a different event from the experimental dataset and the annotated true frame for this video, where the core event occurs. The prompt is illustrated in LABEL:lst:os_prompt. We evaluate the models on each test video in a leave-one-out approach, and average the accuracy for that video across all trials.
@@
The grounding procedure precedes the main test task. The model is instructed to describe the contents of the provided image sequence, and to repeat a particular piece of information from the instructing prompt. The motivation behind this procedure came from a preliminary experiment that was conducted with the online User-Interface model, in which the model had better understanding of the instructing prompt and more attention to the image contents, when explicitly was asked about them. The prompt used for the grounding condition is shown in LABEL:lst:gnd_prompt.
Listing 3: A grounding prompt to verify model’s understanding of both verbal and visual contents of a provided example.
@@
Listing 4: An instructing prompt for detailed reasoning.
@@
The reasoning procedure is combined with the main task, by instructing the model to describe in detail the reasoning behind its answer to the main frame detection task. The motivation for this experimental condition comes from recent studies showing that models’ performance increases when they are explicitly instructed to provide a step-by-step description, and thus may extract and combine more relevant information in the final answer [wei_cot_prompting2022].
@@ 5 Results
(a) Qwen-2.5VL-72B
(b) GPT-4o
S-Figure 3: Mean accuracy vs. detection error tolerance. A correct detection of the models represents a predicted frame within the allowed error tolerance, where an error tolerance of zero means the exact true frame was predicted. Results of Qwen-2.5VL-72B (a) and GPT-4o (b) are shown for the difference ICL regimes under the ”with reasoning” condition. Note, that the length of all videos in the experimental dataset is 10 frames.
@@
The main task for the models was to detect the frame where a test event (’contact’ or ’release) occurs. A correct detection is considered within an allowed error tolerance $e_{\tau}$. We considered an error tolerance in the range: $e_{\tau}\in[0,5]$. For example, $e_{\tau}=0$ means the model predicted the true (human) labeled frame where the event occurs, and $e_{\tau}=1$ means that the predicted frame can be up to a frame off the true frame. Throughout this section, the models’ performance is measured as the mean detection accuracy (with standard error) across all test events (99) in the experimental dataset.
@@
S-Table 2: Reasoning effect on LMMs’ performance. Models are instructed to describe step-by-step the reasoning behind their prediction of the frame where the interaction occurs. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact or two frames off the true frame). Results are reported as per 3 ICL regimes (ZS, OS, TS) for the models: (i) Qwen-2.5VL-72B and (ii) GPT-4o.
ICL Reasoning Mean Accuracy (%)
Exact 2-off
Qwen GPT Qwen GPT
ZS W/O $7\pm 2$ $10\pm 2$ $52\pm 4$ $53\pm 4$
W $\textbf{11}\pm 2$ $9\pm 2$ $\textbf{57}\pm 4$ $55\pm 4$
OS W/O $6\pm 1$ $12\pm 2$ $53\pm 3$ $63\pm 3$
W $\textbf{12}\pm 2$ $11\pm 1$ $\textbf{61}\pm 3$ $62\pm 2$
TS W/O $10\pm 2$ $15\pm 2$ $60\pm 3$ $67\pm 3$
W $\textbf{14}\pm 2$ $13\pm 1$ $\textbf{66}\pm 3$ $66\pm 2$
Fig. 3 depicts the models’ performance as a function of the allowed error tolerance for the models Qwen-2.5VL-72B [Qwen2.5-VL] and GPT-4o [openai2024gpt4o] (other models yielded similar results, see Supplementary). The results shown are under the ”WITH” Reasoning condition (see Sec. 4). The charts show how the very poor performance measured for zero tolerance (TS: 14% and 13% for Qwen and ChatGPT, respectively), increases with a more permitting tolerance, reaching more than 60% for both models. The charts also show a minor increase in performance among the ICL regimes, where TS performs mostly the best and ZS the lowest, as was demonstrated already in earlier studies [dong_icl_survey2024].
Tab. 2 reports the mean accuracy for $e_{\tau}=0$ and $e_{\tau}=2$ (equivalent to Top-1 and Top-5 results) as per each of the three ICL regimes (ZS, OS, TS) and under the Reasoning condition (w or w/o reasoning, see Sec. 4). Similarly, Tab. 3 lists the mean accuracy under the Grounding condition (see Supplementary for more detailed results).
The results in Tab. 2 indicate that instructing the models to provide reasoning for their answers usually increases their performance. For the Qwen-2.5VL model the performance increase is significant, while for GPT-4o it is not (with respect to the standard error). We performed additional ablation experiment, inspired by recent work on boosting LLMs reasoning with verification feedback and similar to reinforcement learning approach [wei_cot_prompting2022, yao_tot2023, guo_deepseekr12025], but this did not improve the models’ performance on the detection task (see Supplementary). Surprisingly, the results in Tab. 3 indicate that in contrast to the Reasoning condition, instructing the model to provide grounding information from the contents of the instruction prompt and video, yields mostly a decrease in the models’ performance. We further discuss this in Sec. 6.

### v2

# Action Without Interaction: Probing the Physical Foundations of Video LMMs via Contact-Release Detection

@@ Abstract
Large multi-modal models (LMMs) show increasing performance in realistic visual tasks for images and, more recently, for videos. For example, given a video sequence, such models are able to describe in detail objects, the surroundings and dynamic actions. In this study, we explored the extent to which these models ground their semantic understanding in the actual visual input. Specifically, given sequences of hands interacting with objects, we asked models when and where the interaction begins or ends. For this purpose, we introduce a first of its kind, large-scale dataset with more than 20K annotated interactions on videos from the Something-Something-V2 dataset. 250 AMTurk human annotators labeled core interaction events, particularly when and where objects and agents become attached (‘contact’) or detached (‘release’). We asked SoTA LMMs, including GPT, Gemini and Qwen to locate these events in short videos, each with a single event. The results show that while models reliably name target objects and identify actions, they exhibit a form of ‘shortcut learning’ where semantic success masks a failure in physical grounding. Specifically, they consistently fail to identify the frame where the interaction begins or ends and poorly localize the physical event within the scene. This disconnect suggests that while LMMs excel at System 1 intuitive pattern recognition (naming the action and objects), they lack the System 2 cognitive foundations required to reason about physical primitives like ‘contact’ and ‘release’, hence truly ground dynamic scenes in physical reality.
@@ 1 Introduction
Discovering and understanding actions and interactions are fundamental cognitive capabilities of humans and other intelligent beings, necessary in interpreting and planning dynamic events between objects and agents in the surrounding environment [22]. Infants develop early sensitivity to spatiotemporal continuity in simple events, for example, the physical contact between an agent and a target object (see Fig. 1).
Figure 1: Events of physical contact or release. Each row presents 2 frames before and after a true physical contact or release event (marked with a green box) during a hand-object interaction. Humans are sensitive to the gap between the hand and the target object even one frame before a contact or after a release. Evidence of a physical contact is crucial to identify an interaction over a simple occlusion, or to predict the outcome of an interaction at it’s termination. These examples are from our annotated dataset, extracted from SSv2.
In developmental psychology, the perception of physical contact and release are considered ‘primitives’ of causal reasoning—foundational building blocks that allow humans to move beyond surface correlations to an understanding of agency and physical laws. This sensitivity to the causality of perception guides infants at a very young age, in learning to detect and interpret interactions between objects and agents, including launching, entraining and expulsion events [18, 14, 23, 2].
@@
Computationally, recent vision models showed increasing performance in recognizing actions and interactions in realistic video sequences [19, 9, 28]. Some models include special architectural designs to improve on the internal representation learning [38], while others use a common architecture, but train on very large unlabeled video datasets utilizing self-supervised learning (SSL) paradigms [30, 16, 5]. The introduction of large multi-modal models (LMMs), allowed the combination of semantic information from large language models (LLMs) and foundational visual representations, thus allowing to generalize to unseen videos and actions without explicit training [28].
@@
Figure 2: Collecting human annotations for interactions using Amazon Mechanical Turk platform. Human subjects were asked to annotate core interaction events in videos from SSv2 dataset [10]. Shown here are example annotations for ‘contact’ and ‘release’ events, where the target object (white candle) comes in contact with a hand (left) and a surface (middle), or is detached from the hand (right). The annotations include the event type, the kind of agent-object pair and the spatiotemporal location of the event (frame and image coordinates).
Despite the increasing success of models in generalizing to more complex tasks without explicit training, recent studies revealed fundamental limitations in the models’ ability to reason about the performed tasks and develop human-like generalizable cognitive understanding [4, 24, 13]. In this study, we explore whether the enhanced performance of LMMs in video action recognition, reflects an improved cognitive understanding (System 2), or merely a superficial ‘story telling’ ability about detected objects in proximity of hands (System 1)?
@@
For this purpose we introduce a large-scale dataset – The Contact-Release Interaction Dataset (CRID). This first of its kind dataset consists of more than 20K annotated interactions, based on $10,000$ action videos from the “Something-Something v.2” (SSv2) dataset [10]. Using the Amazon Mechanical Turk (AMTurk) crowd sourcing platform, we conducted a survey, in which 250 human annotators labeled $24,222$ core interaction events, including: (i) the type of agent acting upon the target object (e.g., a hand or another object), (ii) the type of core interaction event (‘contact’–attachment between a target object and an agent, or ‘release’–detachment of the object from the agent), (iii) the spatiotemporal location of the event (frame number and image coordinates). Fig. 2 depicts the annotation setting (see details in Sec. 3).
@@
Based on these annotations, we conducted a series of experiments to evaluate the ability of current LMMs to detect the spatiotemporal location of core interaction events in real-world video sequences. The experiments were conducted under several In-Context-Learning (ICL) regimes, while applying two modifying conditions on the models’ prompts – Reasoning and Grounding – inspired by earlier studies [25, 32, 12]. We applied the evaluation scheme (see Fig. 3) to five SoTA LMMs including, OpenAI’s GPT-5.2 and GPT-4o models [26, 20], Google’s Gemini-Pro-3 and Gemini-2.5-Flash models [7, 6] and an open source version of Alibaba’s Qwen-2.5VL-72B [1].
@@
  • Introduce CRID - a large scale dataset, based on 10K interaction videos from SSv2 dataset, with more than 20K first of their kind human annotations of core interaction events (‘Contact’ and ‘Release’). The annotations include details on the event and agent types, as well as the spatiotemporal locations of the events.
  • A set of prompting experiments under several ICL regimes and modifying Reasoning and Grounding conditions.
  • A discussion around the striking ‘grounding gap’ revealed in our results: models can identify a high-level action label through statistical association, but remain ‘blind’ to the underlying physical events that define that action.
@@ 2 Related Work
Video understanding was thoroughly researched in the past due to its high value to the advancement in the domain of AI. Unlike action recognition in video, such as identifying people jumping, playing tennis etc., video understanding is a more complex task, often requiring a high level of generalization. Recent studies in this area of research, such as Maaz et al. [17], showed that a ChatGPT agent can successfully answer complex questions when prompted with image data together with the verbal question. Wu et al. [33] used segmentation masks, which the model provided in a grounding step, to further improve the model’s understanding of the input images. Shao et al. [25] showed that by asking the model to produce a Chain-of-Thought of a general task related to the image of the main task, improves significantly the final answer of the LMM. Tian and Wu [29] also used a prompt tuning technique to improve action recognition performance of an LMM agent, but their method requires training of additional adapter models that embed the prompt, actions and the image to be tokenized and sent to the LMM. Chen et al. [3] utilize reinforcement learning to improve LMM grounding via iterative ROI extraction and prediction-based feedback.
While Qi et al. [21] reveal that image-based LMMs often prioritize textual priors over visual evidence, their analysis remains limited to static scenes. Our work extends this critique to the temporal domain, demonstrating that even when models produce coherent action narratives, they fail to ground them in the physical primitives—such as contact and release—that define dynamic interactions.
Finally, several studies focused on changes that a manipulated object undergoes during long interaction videos (e.g., paper folding, fruit peeling, etc.), utilizing existing benchmarks for object affordance and object manipulation [27, 34, 31, 36]. The videos are often at low resolution and include coarse categories of background and actions, rather than the begin or end of a physical contact. In contrast, our evaluation is focused on the initial moment of contact or release in an interaction, regardless of the kind of target object or the kind of manipulation, which are the core of the object affordance studies. We believe that the generic ability to detect physical contact is fundamental for learning about object affordance and manipulations in a self-supervised manner.
@@ 3 Dataset
Figure 3: A schematic flow chart of the experiments under the different In-Context-Learning (ICL) regimes (i.e., ZS, OS, TS) and modulating conditions. The blocks represent different components of intermediate procedures. Each row represents an experiment using a particular ICL regime and condition (the experiment flow is directed left to right). The CNTX block indicates an introductory prompt about the agent. The EXMP block represents a prompt of an example, including the task instruction, an input video and the correct response for this example. The RSN block indicates a prompt instructing the model to include in the response a step-by-step description of the reasoning behind the predicted answer. The GRND block represents a prompt instructing the model to describe the content of the input video and the instructing prompt. In this block, the model provides an intermediate response, prior to the main task. The TST block indicates the prompt of the main test task, including the instruction and test video (see Sec. 4 for more details).
Overview.
In this section we describe the dataset used in our experiments. We explore models’ understanding of core interaction elements, for example the moment a hand picks up a target object and the spatial location where the physical contact occurs in the scene. We introduce a new large scale dataset – The Contact-Release Interaction Dataset (CRID). This first of its kind dataset consists of more than 20K annotated interactions on more than 10K action videos from the “Something-Something-v2” (SSv2) dataset, which is a collection of smartphones’ videos depicting humans performing various everyday interactions in natural settings [10]. The SSv2 labels include generic action templates, for example, “putting something into something”, and the object names in the template placeholders (“something”) for each video.
Table 1: The Contact-Release Interaction Dataset (CRID). New annotations of core interaction events for videos extracted from the SSv2 dataset [10].
@@
  ‘Contact’ events   $13816$
  ‘Release’ events   $10406$
  ‘Hand-object’ interactions   $12550$
  ‘Object-object’ interactions   $5653$
  ‘Object-surface’ interactions   $6019$
Human annotations.
In this work we used the AMTurk online platform to conduct a survey, in which 250 human annotators labeled more than 20K core interaction events, including: (i) the type of agent acting upon the target object (e.g., a hand or another object), (ii) the type of core interaction event, i.e, ‘contact’ or ‘release’, (iii) the spatiotemporal location of the event, i.e., the frame number and image coordinates where the physical event occurs (see Fig. 1 and Fig. 2). The survey and the annotations collection procedure were approved by the institutional review board of the Weizmann Institute of Science, Rehovot, Israel. All human subjects gave informed consent before participating in the survey. Details about the new annotated dataset, including the videos and the labels are summarized in Tab. 1. Due to time and budget constraints, we could not hire multiple annotators for each of the 10K videos. However, we evaluated the inter-rater agreement between randomly selected annotators, by measuring the Intraclass Correlation Coefficient (ICC) between 3 random annotators on a randomly selected subset of videos. The results indicate a high agreement level on the frame annotation (ICC of 0.95), and a medium agreement level on the spatial location where the events occur (ICC of 0.73 and 0.39 for the exact x and y coordinates, respectfully, and an IoU of 0.57 for a $\times120$ bounding box centered around the labeled location).
Open access availability.
The dataset is publicly available at ssv2-contact-release-interaction-dataset. Despite concerns about benchmark data leaking into LMM pre-training, recent research (e.g., [11]) indicates that even top-tier SSL visual encoders struggle with physical interactions and contact, even after fine-tuning. Consequently, we provide open access to our annotations. This allows future models to leverage these scarce examples alongside their inherent reasoning and semantic knowledge (similar to humans) to bridge current limitations.
@@ 4 Experimental Design
In this section we describe in more detail the evaluation experiments conducted in the course of this study. We considered three ICL regimes: zero-shot, one-shot and two-shot [8].
As we are interested in models’ ability to understand core interaction events in videos, we focused on the detection of the frame in the video, where a core interaction event (‘contact’ or ‘release’) occurs. Since the original SSv2 videos often contain several simultaneous interaction events, we manually extracted 3 temporally isolated events from a subset of 33 videos, resulting with 99 short sequences, 10 frames each, consisting of a single core event (see supplementary for details). As a control, we extracted from each video, an additional short sequence (a ‘non-event’), which depicts objects and hands in a dynamic scene without a physical contact/release event.
To better understand the models behavior, we employed two explainability methods, that were used in previous studies with LMMs, as modulating conditions: (i) Grounding, (ii) Reasoning. We elaborate about these conditions later in this section. In our experiments we evaluated their relative influence on the models performance. These prompting conditions do not modify the models’ internal reasoning parameters. Fig. 3 presents the different experimental settings, including the ICL regimes and modulating conditions.
@@
In the OS regime, models are provided with a single example of the main task prior to performing the task on the test video sequence. The example, includes a video with a different event from the experimental dataset and the annotated true frame for this video, where the core event occurs. The prompt is provided in LABEL:lst:os_prompt. We evaluate the models on each test video in a leave-one-out approach, and average the accuracy for that video across all trials.
@@
The grounding procedure precedes the main test task. The model is instructed to describe the contents of the provided image sequence, and to repeat a particular piece of information from the instructing prompt. The motivation behind this procedure came from a preliminary experiment that was conducted with the online User-Interface of ChatGPT-4o, in which the model had better understanding of the instructing prompt and more attention to the image contents, when explicitly was asked about them. The prompt used for the grounding condition is shown in LABEL:lst:gnd_prompt.
Listing 3: A grounding prompt in the OS experiment to verify model’s understanding of both verbal and visual contents of a provided example.
@@
Listing 4: An instructing prompt for step-by-step reasoning.
@@
The reasoning procedure is combined with the main task, by instructing the model to describe in detail the reasoning behind its answer to the main frame detection task (see prompt in LABEL:lst:rsn_prompt). This prompt does not modify the internal reasoning parameters of the model. The motivation for this experimental condition comes from recent studies showing that models’ performance increases when they are explicitly instructed to provide a step-by-step description, and thus may extract and combine more relevant information in the final answer [32]. Fig. 4 shows an example of a false prediction together with the model’s step-by-step chain of thought.
Figure 4: Example false prediction. The model provides the chain-of-thought under the “WITH” Reasoning condition. The example shows that the reasoning seems logical and realistic, but the grounding of the physical release event in the video frames is very loose. The orange box marks the true frame. The red box marks the false prediction.
@@ 5 Results
The main task for the models was to detect the frame where a test event (’contact’ or ’release) occurs. A correct detection is considered within an allowed error tolerance $e_{\tau}$. For example, $e_{\tau}=0$ means the model predicted the exact true (human) labeled frame where the event occurs, and $e_{\tau}=1$ means that the predicted frame can be up to one-frame off (before or after) the true frame.
@@
Table 2: Reasoning effect on LMMs’ performance. Models are instructed to describe step-by-step the reasoning behind their prediction of the frame where the interaction occurs. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact or one frame off the true frame). Results are reported as per 3 ICL regimes (ZS, OS, TS) for the models: (i) Gemini-2.5-Flash and (ii) GPT-5.2.
ICL RSN Mean Accuracy Percentage (SD)
Exact 1-off
Gemini GPT Gemini GPT
ZS W/O $8.08$ $(1.94)$ $9.60$ $(2.10)$ $30.81$ $(3.29)$ 35.35 $(3.41)$
W $9.09$ $(2.05)$ 12.12 $(2.33)$ $29.80$ $(3.26)$ $33.84$ $(3.37)$
OS W/O 19.87 $(2.84)$ $18.15$ $(2.29)$ 47.15 $(3.55)$ $42.96$ $(3.19)$
W $16.74$ $(2.65)$ $15.93$ $(2.02)$ $43.71$ $(3.53)$ $39.43$ $(3.05)$
TS W/O 20.71 $(20.71)$ $18.03$ $(2.73)$ 49.07 $(3.55)$ $44.31$ $(3.53)$
W $17.98$ $(2.73)$ $17.12$ $(2.68)$ $47.53$ $(3.55)$ $40.82$ $(3.49)$
As demonstrated in Fig. 1, it is often evident that there is no physical contact between the hand and target object already one frame after a release or before a contact event. Throughout this section, the models’ performance is measured as the mean detection accuracy (with standard error) across all test events (99) in the experimental dataset.